OpenStudy (truwhovian11):

SOS WILL MEDAL AND FAN WHOEVER HELPS ME FIRST Convert the decimal to a percent. 0.007 A. 0.007% B. 0.07% C. 0.7% D. 7%

OpenStudy (truwhovian11):

@hperez777 @jabez177 @LegendarySadist

OpenStudy (truwhovian11):

@Comrad

OpenStudy (truwhovian11):

hello @wilsondanielle

OpenStudy (truwhovian11):

r u there

OpenStudy (wilsondanielle):

To convert from a decimal to a percent, multiply by 100. Thus, 0.007*100 will be the percent

OpenStudy (truwhovian11):

k if no one will help ill check my self

OpenStudy (truwhovian11):

ok thanks

OpenStudy (truwhovian11):

can u help on somemore

OpenStudy (wilsondanielle):

yes

OpenStudy (truwhovian11):

thx

OpenStudy (truwhovian11):

Convert the percent to a decimal. 12.5% A. 0.00125 B. 0.125 C. 1.25 D. 12.5

OpenStudy (truwhovian11):

Which percent is equal to the given fraction? 3/5 A. 3.5% B. 30% C. 35% D. 60%

OpenStudy (wilsondanielle):

This will be the opposite - divide by 100 to get the decimal. This is because 100% is equal to 1 - so to convert from decimal (which is between 0 and 1 in most cases UNLESS the total percent is over 100) multiply by 100, and to convert from percent divide by 100 (in most cases this will yield a decimal between 0 and 1, again unless the percent is over 100).

OpenStudy (truwhovian11):

for number 1 i think a or b

OpenStudy (wilsondanielle):

For number 1 : 0.007*100% = 0.7% = C

OpenStudy (truwhovian11):

oh ok so number 2 a or b ???

OpenStudy (wilsondanielle):

For number 2 : 12.5%/100% = 0.125 = B Remember - the decimal should be between 0 and 1 if the percent is between 0 and 100.

OpenStudy (truwhovian11):

um B

OpenStudy (truwhovian11):

i think

OpenStudy (wilsondanielle):

What about 3? Hint : start by converting the fraction into a decimal

OpenStudy (truwhovian11):

@wilsondanielle

OpenStudy (truwhovian11):

how do i do that again sorry

OpenStudy (truwhovian11):

@wilsondanielle

OpenStudy (truwhovian11):

@ayeedomo

OpenStudy (wilsondanielle):

3/5 can be punched into your calculator for the decimal. (3/5 = 0.6) Alternatively, you can go straight for the percent by dividing 100 into 5 portions and cross multiplying (100/5 = 20; therefore each fifth is equal to 20%. 3*20% = 60%)

OpenStudy (wilsondanielle):

Sorry for the slow replies - I'm at work and the computer is from the 90's.

OpenStudy (truwhovian11):

lol my connection is slo 2 so we both have computer problems

OpenStudy (truwhovian11):

@Tal_Tal

OpenStudy (truwhovian11):

thx wish i could medal u twice but i still ned help on a question @wilsondanielle ill make another question

OpenStudy (wilsondanielle):

ok

jabez177 (jabez177):

Nice job guiding the user to the answer. :)